Please help me to solve this equation
5= (4x+1)
2 (3x+1) (3x+1) = 0
7 (2x - 1)
2 (3x +5)
8x + 3 = -4x

Answer by Alan3354(69443)

5= (4x+1)
Subtract
4 = 4x
Divide by 4
x = 1
======
2 (3x+1) (3x+1) = 0
(3x+1)*(3x+1) = 0
3x+1 = 0
3x = -1
x = -1/3
==============
7 (2x - 1)
No equal sign - not an equation.
2 (3x +5)
No equal sign - not an equation.
=======
8x + 3 = -4x
Add 4x
Subtract 3
etc